# Lie Detector: Unified Backdoor Detection via Cross-Examination Framework

### v1

@@ 1 Introduction
Deep learning models have grown exponentially in size in recent years, outstripping the computational resources available to many small and medium-sized institutions. Consequently, these institutions often rely on third-party cloud providers for model training. Although these providers are considered “semi-honest” in that they ostensibly adhere to prescribed protocols, they may still covertly manipulate data or models. This scenario can give rise to a significant backdoor threat, where hidden triggers are embedded during training, enabling the model to function normally under most conditions but exhibit malicious behavior when specific triggers are activated [10, 13, 2, 19, 22].
Current backdoor detection methods frequently rely on model behavior and statistical analyses (e.g., gradient-based detection, posterior analysis) [30, 24, 20, 9, 31, 25, 37]. However, such approaches tend to be highly sensitive to variations in optimization objectives, loss functions, and feature representations across different learning paradigms. This limitation constrains their ability to generalize across diverse architectures and attack strategies, posing serious challenges for maintaining user model security in a semi-honest setting.
@@
We extensively evaluate the effectiveness of Lie Detector across supervised, semi-supervised, and autoregressive learning paradigms. The results show that our method significantly outperforms state-of-the-art backdoor detection approaches, with relative improvements of +1.6%, +5.4%, and +11.9%, respectively. In addition, Lie Detector demonstrates high stability under varying random seeds, underscoring its robustness. We anticipate that this research will encourage broader adoption of secure training practices in third-party services, thereby strengthening security guarantees for deep learning models. Our contributions are
@@ 5.1 Implementation Details
Attacks and Defenses. We consider backdoor attacks across different paradigms, including BadNets [10], Blended [2], ISSBA [19], WaNet [26], and Low-Frequency [34] for supervised learning. For self-supervised and autoregressive learning, we adapt these attacks and further evaluate BadCLIP [22], BadEncoder [13], TrojanVLM [21], and Shadowcast [33]. We employ advanced defenses, including NC [30], ABS [24], NAD [20], TED [25], MM-BD [31], DECREE [9], and SEER [37]. Some methods, such as TED and MM-BD, are extended to multiple paradigms. Unless otherwise specified, all attack methods use a 10% poisoning rate. All evaluations are conducted using the semi-honest environment, with detailed settings and evaluation metrics provided in LABEL:app:attack and LABEL:app:defense.
@@ 5.4 Ablation Study
Similarity metrics selection. We select four existing similarity metrics for comparative testing, including CKA, CCA (Canonical Correlation Analysis), SVCCA (Singular Vector Canonical Correlation Analysis) and COS (Cosine Similarity) [16]. We adaptively replace the aforementioned different similarity metrics for backdoor detection, and the results are shown in LABEL:fig:f1_four. The results indicate that CKA outperforms other similarity metrics across different learning paradigms, especially on the LLaVA model, which also indirectly demonstrates the architecture-agnostic nature of the CKA metric.

### v2

@@ 1 Introduction
Deep learning models have grown exponentially in size in recent years, outstripping the computational resources available to many small and medium-sized institutions. Consequently, these institutions often rely on third-party cloud providers for model training. Although these providers are considered “semi-honest” in that they ostensibly adhere to prescribed protocols, they may still covertly manipulate data or models. This scenario can give rise to a significant backdoor threat, where hidden triggers are embedded during training, enabling the model to function normally under most conditions but exhibit malicious behavior when specific triggers are activated [11, 31, 2, 20, 24, 22, 29, 26, 49].
Current backdoor detection methods frequently rely on model behavior and statistical analyses (e.g., gradient-based detection, posterior analysis) [38, 32, 21, 10, 39, 33, 48, 40]. However, such approaches tend to be highly sensitive to variations in optimization objectives, loss functions, and feature representations across different learning paradigms [4]. This limitation constrains their ability to generalize across diverse architectures and attack strategies [30, 45, 42, 25], posing serious challenges for maintaining user model security in a semi-honest setting.
@@
We extensively evaluate the effectiveness of Lie Detector across supervised, semi-supervised, and autoregressive learning paradigms. The results show that our method significantly outperforms state-of-the-art backdoor detection approaches, with relative improvements of +5.4%, +1.6%, and +11.9%, respectively. In addition, Lie Detector demonstrates high stability under varying random seeds, underscoring its robustness. We anticipate that this research will encourage broader adoption of secure training practices in third-party services, thereby strengthening security guarantees for deep learning models. Our contributions are
@@ 5.1 Implementation Details
Attacks and Defenses. We consider backdoor attacks across different paradigms, including BadNets [11], Blended [2], ISSBA [20], WaNet [34], and Low-Frequency [44] for supervised learning. For self-supervised and autoregressive learning, we adapt these attacks and further evaluate BadCLIP [27], BadEncoder [14], TrojanVLM [23], and Shadowcast [43]. We employ advanced defenses, including NC [38], ABS [32], NAD [21], TED [33], MM-BD [39], DECREE [10], and SEER [48]. Some methods, such as TED and MM-BD, are extended to multiple paradigms. Unless otherwise specified, all attack methods use a 10% poisoning rate. All evaluations are conducted using the semi-honest environment, with detailed settings and evaluation metrics provided in Appendices C.1 and C.2.
@@ 5.4 Ablation Study
Similarity metrics selection. We select four existing similarity metrics for comparative testing, including CKA, CCA (Canonical Correlation Analysis), SVCCA (Singular Vector Canonical Correlation Analysis) and COS (Cosine Similarity) [17]. We adaptively replace the aforementioned different similarity metrics for backdoor detection. The results indicate that CKA outperforms other similarity metrics across different learning paradigms, especially on the LLaVA model, which also indirectly demonstrates the architecture-agnostic nature of the CKA metric.